OpenStudy (anonymous):

Trigonometry find the distance between (4/5, -2/3) and (-8/5. -6). also the MIDPOINT and the SLOPE. Trigonometry find the distance between (4/5, -2/3) and (-8/5. -6). also the MIDPOINT and the SLOPE. @Mathematics

OpenStudy (lalaly):

\[\underline{distance}=\sqrt{(y_2-y_1)^2 +(x_2 -x_1)^2}\]\[\underline{midpoint}= (\frac{x_1+x_2}{2} ,\frac{y_2+y_1}{2})\]\[\underline{slope} = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

can you help me in solving that?

OpenStudy (lalaly):

try doing it, and i'll check ur answers for u if u want... i gave u the solution, u just need to plug in the pointts and ull get the answers

OpenStudy (anonymous):

ok:))

OpenStudy (anonymous):

in our book its x first b4 y..

OpenStudy (anonymous):

in solving the distance

OpenStudy (lalaly):

it doesnt matter,

OpenStudy (anonymous):

d= \[\sqrt{7896/225}\]

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

is it right?